# Masked Diffusion models are Secretly Time-Agnostic Masked Models and Exploit Inaccurate Categorical Sampling

### v1

@@ Abstract
Masked diffusion models (MDMs) have emerged as a popular research topic for generative modeling of discrete data, thanks to their superior performance over other discrete diffusion models, and are rivaling the auto-regressive models (ARMs) for language modeling tasks. The recent effort in simplifying the masked diffusion framework further leads to alignment with continuous-space diffusion models and more principled training and sampling recipes. In this paper, however, we reveal that both training and sampling of MDMs are theoretically free from the time variable, arguably the key signature of diffusion models, and are instead equivalent to masked models. The connection on the sampling aspect is drawn by our proposed first-hitting sampler (FHS). Specifically, we show that the FHS is theoretically equivalent to MDMs’ original generation process while significantly alleviating the time-consuming categorical sampling and achieving a 20$\times$ speedup. In addition, our investigation challenges previous claims that MDMs can surpass ARMs in generative perplexity. We identify, for the first time, an underlying numerical issue, even with the 32-bit floating-point precision, which results in inaccurate categorical sampling. We show that the numerical issue lowers the effective temperature both theoretically and empirically, leading to unfair assessments of MDMs’ generation results in the previous literature.
@@ 1 Introduction
In this paper, we dive deeper into MDMs and their connection to diffusion models and masked models. We highlight our key findings: (1) MDMs, in both training and sampling, are essentially time-agnostic masked models (or order-agnostic auto-regressive models), enjoying 20$\times$ faster sampling and diverging from the design choices of diffusion models. This also justifies the theoretical foundation of masked models as they are equivalent and simpler formulations of the more principled MDMs. (2) We challenge previous claims that MDMs can surpass ARMs in generative perplexity by identifying a hidden but critical numerical issue that renders previous evaluations unfair. After correcting it, we find that MDMs significantly lag behind ARMs in generative perplexity.
@@ 3.1 Reformulating the ELBO with the Number of Masked Tokens
From Eqn. (7), the sequence NELBO can be expressed discretely with the time-agnostic network $\bar{\bm{\mu}}_{\theta}(\bm{x})$. Therefore, Eqn. (7) can serve as a NELBO of masked models in a straightforward approach: uniformly choose the number of masked tokens $n$ from $\{1,\dots,L\}$, uniformly mask $n$ random tokens in $\bm{x}_{0}$ to obtain $\bm{x}_{n}$, and compute the average cross-entropy loss of $\bar{\bm{\mu}}_{\theta}(\bm{x})$ on these $n$ positions. The weighting $\frac{1}{n}$ in this NELBO resembles the likelihood weighting in diffusion models (Song et al., 2021b; Kingma et al., 2021; Lu et al., 2022a; Zheng et al., 2023b), facilitating maximum likelihood training of masked models. Note that early works on order-agnostic auto-regressive models (Uria et al., 2014; Hoogeboom et al., 2021a) already reveal this weighting from a different perspective. While in the context of masked models, there are few discussions on the ELBO. Discussions on related work are placed in Appendix B.
@@ 4.1 Inefficiency of Current Sampling
When using the caching strategy in batched sampling, the network output can only be reused directly when all the sequences in the batch remain unchanged after a sampling step. Suppose the batch size is $B$, and the default linear noise schedule $\alpha_{t}=1-t$ as well as uniform timesteps $t_{k}=\frac{k}{N}$ is used. The expected NFE under the caching strategy can be derived as $N(1-(1-\frac{1}{N})^{BL})$ (proof in Appendix E) , similar to the $B=1$ case in Ou et al. (2024). As $\lim_{N\rightarrow\infty}N(1-(1-\frac{1}{N})^{BL})=BL$, the NFE is no longer upper bounded by the sequence length but scales with the batch size (Figure 3b).
@@ 4.2 First-Hitting Sampler
When the number of sampling steps $N\rightarrow\infty$ and the maximum step size $\max_{1\leq i\leq N}|t_{i}-t_{i-1}|\rightarrow 0$, Eqn. (10) tends to an infinitesimal jump. In this case, the reverse sampling process becomes a continuous-time Markov chain (or Markov process), where each mask token is unmasked at some moment according to the network prediction. Our key insight involves three folds: (1) Whether a mask token will transit or not during a time interval $[s,t]$ is independent of the network. The network output only determines which token is the transition target given the condition that the transition happens. (2) The transition probability $\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}$ is equal for masked tokens at different positions. Therefore, each mask token has the same probability of being first unmasked. (3) The first-hitting time, which denotes the first moment any of the remaining masked tokens is unmasked, can be analytically sampled according to the following proposition:
@@ 5 Are MDMs Better than ARMs? A Critical Fault in Low-Precision Gumbel-Based Categorical Sampling
Before we proceed to verify the effectiveness of our proposed first-hitting sampler, we have to point out a critical fault in MDMs’ original sampling. As suggested by previous works (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024), MDMs seem to surpass ARMs with a sufficient number of sampling steps when measured by the generative perplexity (Gen PPL) [^3] The evaluation metrics used in this paper are introduced in Appendix H.1., as shown in Figure 7a. However, in this section, we identify for the first time a hidden implementation issue existing in previous codebases that leads to this false conclusion.
@@ 5.1 Low Token Diversity under Numerous Sampling Steps
However, when we check the generated content, we discover that the quality is compromised by low token diversity (an extreme case is shown in Figure 6). We further quantify this phenomenon by measuring the sentence entropy (Figure 7b). With the original sampler, the generative perplexity of MDMs surpasses ARMs at around 2k steps, but the entropy is always lower and keeps decreasing.
@@ 5.2 Identifying the Numerical Precision Problem
This low generation quality is unexpected, as theory suggests that approximation error should decrease with more sampling steps. Consequently, samples generated over numerous steps should reflect the true performance of MDMs. We therefore consider this an implementation issue and investigate further to identify the root cause.
@@
Note that auto-regressive LLMs like Llama (Touvron et al., 2023) and Mistral (Jiang et al., 2023) use torch.multinomial for categorical sampling, which is also implemented with the Gumbel-max trick but fixed as 64-bit precision in the low-level C++ code of PyTorch. Besides, we find the truncation less influential in ARMs.
@@ 7 Conclusion
In this work, we advance our understanding of masked diffusion models (MDMs) by revealing their theoretical equivalence to masked models and addressing a critical implementation issue that compromised the fairness of previous evaluations of MDMs’ generative performance. Our investigation challenges earlier claims that MDMs can surpass ARMs in text generation. Despite our negative findings, we acknowledge that our text-based experiments may inherently favor ARMs, as text naturally follows a left-to-right order that ARMs are better suited to model. We believe that MDMs are potentially well-suited for applications where the data’s order-agnostic nature is a key prior.
@@ I.1.2 Failed Training Attempts
We visualize the training curves and optimized importance weights at around 100k iterations in Figure 12c. Unfortunately, while the adaptive IS technique reduces the variance to a level comparable to the low-discrepancy sampler, it also results in degraded performance. We hypothesize that the dynamical updated $p_{t}$ may make the loss estimator biased.

### v2

@@ Abstract
Masked diffusion models (MDMs) have emerged as a popular research topic for generative modeling of discrete data, thanks to their superior performance over other discrete diffusion models, and are rivaling the auto-regressive models (ARMs) for language modeling tasks. The recent effort in simplifying the masked diffusion framework further leads to alignment with continuous-space diffusion models and more principled training and sampling recipes. In this paper, however, we reveal that both training and sampling of MDMs are theoretically free from the time variable, arguably the key signature of diffusion models, and are instead equivalent to masked models. The connection on the sampling aspect is drawn by our proposed first-hitting sampler (FHS). Specifically, we show that the FHS is theoretically equivalent to MDMs’ original generation process while significantly alleviating the time-consuming categorical sampling and achieving a 20$\times$ speedup. In addition, our investigation raises doubts about whether MDMs can truly beat ARMs. We identify, for the first time, an underlying numerical issue, even with the commonly used 32-bit floating-point precision, which results in inaccurate categorical sampling. We show that the numerical issue lowers the effective temperature both theoretically and empirically, and the resulting decrease in token diversity makes previous evaluations, which assess the generation quality solely through the incomplete generative perplexity metric, somewhat unfair.
@@ 1 Introduction
In this paper, we dive deeper into MDMs and their connection to diffusion models and masked models. We highlight our key findings: (1) MDMs, in both training and sampling, are essentially time-agnostic masked models (or order-agnostic auto-regressive models), enjoying 20$\times$ faster sampling and diverging from the design choices of diffusion models. This also justifies the theoretical foundation of masked models as they are equivalent and simpler formulations of the more principled MDMs. (2) We challenge previous claims that MDMs can surpass ARMs by identifying a hidden but critical numerical issue that reduces the token diversity and renders previous evaluations unfair. After correcting it, we find that MDMs significantly lag behind ARMs in generative perplexity.
@@ 3.1 Reformulating the ELBO with the Number of Masked Tokens
From Eqn. (7), the sequence NELBO can be expressed discretely with the time-agnostic network $\bar{\bm{\mu}}_{\theta}(\bm{x})$. Therefore, Eqn. (7) can serve as a NELBO of masked models in a straightforward approach: uniformly choose the number of masked tokens $n$ from $\{1,\dots,L\}$, uniformly mask $n$ random tokens in $\bm{x}_{0}$ to obtain $\bm{x}_{n}$, and compute the average cross-entropy loss of $\bar{\bm{\mu}}_{\theta}(\bm{x})$ on these $n$ positions. The weighting $\frac{1}{n}$ in this NELBO resembles the likelihood weighting in diffusion models (Song et al., 2021b; Kingma et al., 2021; Lu et al., 2022a; Zheng et al., 2023b), facilitating maximum likelihood training of masked models. Note that early works on order-agnostic auto-regressive models (Uria et al., 2014; Hoogeboom et al., 2021a) already reveal this weighting from a different perspective [^1] The relation between ELBOs of order-agnostic ARMs and MDMs was also mentioned in a recent work (Ou et al., 2024), while they only consider an originally time-agnostic network instead of mixture-of-experts.. While in the context of masked models, there are few discussions on the ELBO. Discussions on related work are placed in Appendix B.
@@ 4.1 Inefficiency of Current Sampling
When using the caching strategy in batched sampling, the network output can only be reused directly when all the sequences in the batch remain unchanged after a sampling step [^4] We can reuse only the unchanged part of a batch, but this potentially reduce parallel efficiency.. Suppose the batch size is $B$, and the default linear noise schedule $\alpha_{t}=1-t$ as well as uniform timesteps $t_{k}=\frac{k}{N}$ is used. The expected NFE under the caching strategy can be derived as $N(1-(1-\frac{1}{N})^{BL})$ (proof in Appendix E) , similar to the $B=1$ case in Ou et al. (2024). As $\lim_{N\rightarrow\infty}N(1-(1-\frac{1}{N})^{BL})=BL$, the NFE is no longer upper bounded by the sequence length but scales with the batch size (Figure 3b).
@@ 4.2 First-Hitting Sampler
When the number of sampling steps $N\rightarrow\infty$ and the maximum step size $\max_{1\leq i\leq N}|t_{i}-t_{i-1}|\rightarrow 0$, Eqn. (10) tends to an infinitesimal jump. In this case, the reverse sampling process becomes a continuous-time Markov chain (or Markov process), where each mask token is unmasked at some moment according to the network prediction. Our key insight involves three folds: (1) Whether a mask token will transit or not during a time interval $[s,t]$ is independent of the network. The network output only determines which token is the transition target given the condition that the transition happens. (2) The transition probability $\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}$ is equal for masked tokens at different positions. Therefore, each mask token has the same probability of being first unmasked. (3) The first-hitting time, which denotes the first moment any of the remaining masked tokens is unmasked, can be analytically sampled:
@@ 5 Are MDMs Better than ARMs? A Critical Fault in Low-Precision Gumbel-Based Categorical Sampling
Before we proceed to verify the effectiveness of our proposed first-hitting sampler, we have to point out a critical fault in MDMs’ original sampling. As suggested by previous works (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024), MDMs seem to surpass ARMs with a sufficient number of sampling steps when measured by the generative perplexity (Gen PPL) [^5] The evaluation metrics used in this paper are introduced in Appendix H.1., as shown in Figure 7a. However, in this section, we identify for the first time a hidden implementation issue existing in previous codebases that makes this conclusion doubtful.
@@ 5.1 Low Token Diversity under Numerous Sampling Steps
However, when we check the generated content, we discover that the quality is compromised by low token diversity (an extreme case is shown in Figure 6). We further quantify this phenomenon by measuring the sentence entropy (Figure 7b). With the original sampler, the Gen PPL of MDMs surpasses ARMs at around 2k steps, but the entropy is always lower and keeps decreasing.
@@ 5.2 Identifying the Numerical Precision Problem
This low generation quality is unexpected, as theory suggests that approximation error should decrease with more sampling steps. Consequently, samples generated over numerous steps should reflect the true performance of MDMs. We therefore consider this a hidden implementation issue and investigate further to identify the root cause.
@@
Note that auto-regressive LLMs like Llama (Touvron et al., 2023) and Mistral (Jiang et al., 2023) use torch.multinomial for categorical sampling, which is also implemented with the Gumbel-max trick in the low-level C++ code of PyTorch. In contrast, we find the truncation under 32-bit precision not influential in the token-by-token decoding process of ARMs.
@@ 7 Conclusion
In this work, we advance our understanding of masked diffusion models (MDMs) by revealing their theoretical equivalence to masked models and addressing a hidden implementation issue that compromised the fairness of previous evaluations of MDMs’ generative performance. Our investigation challenges earlier claims that MDMs can surpass ARMs in text generation. Despite our negative findings, we acknowledge that our text-based experiments may inherently favor ARMs, as text naturally follows a left-to-right order that ARMs are better suited to model. We believe that MDMs are potentially well-suited for applications where the data’s order-agnostic nature is a key prior.
@@ I.1.2 Failed Training Attempts
We visualize the training curves and optimized importance weights at around 100k iterations in Figure 12. Unfortunately, while the adaptive IS technique reduces the variance to a level comparable to the low-discrepancy sampler, it also results in degraded performance. We hypothesize that the dynamical updated $p_{t}$ may make the loss estimator biased.